OpenStudy (waheguru):

find the equation of a line if it has the same y intercept as 3x-7y+21=0 snd is parallel to the x axix

OpenStudy (waheguru):

i got 7= 24/7

OpenStudy (waheguru):

i got y = 24/7

OpenStudy (s3a):

y = 0x - 21/3? I did it in my head so I could be wrong.

OpenStudy (anonymous):

y =3

OpenStudy (s3a):

Isn't it only the y-intercept that it takes from the given equation? And the slope from the line y = 0x?

OpenStudy (aroub):

sorry! misread the question :(

OpenStudy (anonymous):

Aroubt did you know about the form of straight line $$ \frac x a +\frac yb =1 $$

OpenStudy (anonymous):

s3a's approach is a shortcut but not good for a beginner IMO

OpenStudy (s3a):

And the answer should be y = 0x +21/3 (not negative)

OpenStudy (s3a):

Basically, waheguru, the number in front of the x is the slope and the number that is alone is the y-intercept of a function. This question is asking to find another function that has the same number without the x near it and the same number in front of the x as the equation y = 0x. Why y = 0x because that's the equation of the x-axis.